# Small Models Struggle to Learn from Strong Reasoners

### v1

@@ Abstract
Yuetai Li Xiang Yue Zhangchen Xu Fengqing Jiang Luyao Niu Bill Yuchen Lin Bhaskar Ramasubramanian Radha Poovendran University of Washington Carnegie Mellon University Western Washington University {yuetaili,zxu9,fqjiang,luyaoniu,byuchenl,rp3}@uw.edu, xyue2@andrew.cmu.edu, ramasub@wwu.edu
@@ 3.2 Long CoT Gap
Llama-3.2-8B | 25.6 | 25.1 | 0.5 | Strong |
Llama-3.2-70B | 57.6 | 53.3 | 4.3 | Strong |
@@ 3.3 Large Teacher CoT Gap
Strong Model CoT | 57.5 | 35.0 | 80.0 | 25.9 | 0.0 | 39.7 | |
Weak Model CoT | 60.3 | 27.5 | 79.5 | 26.4 | 3.3 | 39.4 | |
@@
Strong Model CoT | 47.4 | 25.0 | 71.2 | 16.9 | 3.3 | 32.8 | |
Weak Model CoT | 47.9 | 17.5 | 74.1 | 16.4 | 3.3 | 31.2 | |
@@ A.2 Training Setup
Teacher models generate responses by rejection sampling (Zelikman et al., 2022; Tong et al., 2024; Yue et al., 2023; Singh et al., 2024; Gulcehre et al., 2023; Yuan et al., 2023; Dong et al., 2023). By default, teacher models employ greedy decoding. By combining the math problem instructions with corresponding solutions generated by teacher models, we construct problem-solution pairs to fine-tune student models. We perform pairwise comparisons of solutions generated by different teacher models and filter out problem-solution pairs that are correct for both models to fine-tune student models.

### v2

@@ Abstract
Yuetai Li Xiang Yue Zhangchen Xu Fengqing Jiang Luyao Niu Bill Yuchen Lin Bhaskar Ramasubramanian Radha Poovendran University of Washington Carnegie Mellon University Western Washington University {yuetaili,zxu9,fqjiang,luyaoniu,byuchenl,rp3}@uw.edu, xyue2@andrew.cmu.edu, ramasub@wwu.edu Huggingface: https://huggingface.co/UWNSL Project Page: https://small-model-gap.github.io/
@@ 3.2 Long CoT Gap
Llama-3.1-8B | 25.6 | 25.1 | 0.5 | Strong |
Llama-3.3-70B | 57.6 | 53.3 | 4.3 | Strong |
@@ 3.3 Large Teacher CoT Gap
Large Model CoT | 57.5 | 35.0 | 80.0 | 25.9 | 0.0 | 39.7 | |
Small Model CoT | 60.3 | 27.5 | 79.5 | 26.4 | 3.3 | 39.4 | |
@@
Large Model CoT | 47.4 | 25.0 | 71.2 | 16.9 | 3.3 | 32.8 | |
Small Model CoT | 47.9 | 17.5 | 74.1 | 16.4 | 3.3 | 31.2 | |
@@ A.2 Training Setup
Teacher models generate responses by rejection sampling (Zelikman et al., 2022; Tong et al., 2024; Yue et al., 2023; Singh et al., 2024; Gulcehre et al., 2023; Yuan et al., 2023; Dong et al., 2023). The prompt used to make teacher models generate responses to the math questions is shown as follows. By default, teacher models employ greedy decoding. By combining the math problem instructions with corresponding solutions generated by teacher models, we construct problem-solution pairs to fine-tune student models. We perform pairwise comparisons of solutions generated by different teacher models and filter out problem-solution pairs that are correct for both models to fine-tune student models.